OpenStudy (anonymous):

Rewrite with positive exponents. Assume that even roots are of nonnegative quantities and that all denominators are nonzero. (4xy ^– 2)^– 2

jhonyy9 (jhonyy9):

y^(-2) =1/y^2 so than will be (4x/y^2)^-2 what will be 1 1 y^4 ----------- = ----------- = ------- (4x/y^2)^2 16x^2 16x^2 -------- y^4